Convert 50°c to fahrenheit

1
Expert's answer
2021-02-07T19:06:46-0500

We can convert celsius to Fahrenheit using the following formula:


"T_{^{\circ}F}=T_{^{\circ}C}\cdot \dfrac{9}{5}+32,""T_{^{\circ}F}=50^{\circ}C\cdot \dfrac{9}{5}+32=122^{\circ}F"